OpenStudy (fluttershyk):

Which table shows a proportional relationship between x and y? (tables in comments)

OpenStudy (fluttershyk):

@KamiBug @rebeccaxhawaii

OpenStudy (welshfella):

look for a table where each value of y can be found by multiplying the x value by the same number.

OpenStudy (fluttershyk):

the second one

OpenStudy (welshfella):

if we look at table 1 :- x = 1 y = 5 1*5 = 5 x = 3 y = 12 3*4 = 12 so this is not a proportional relation because we are multiplying x by different numbers

OpenStudy (fluttershyk):

I don't get it none of them do that

OpenStudy (anonymous):

okay it's easy @fluttershyK

OpenStudy (anonymous):

like to learn the shortcut

OpenStudy (welshfella):

Yes one does Check out table 2 Another way to check would be to calculate y/x and see if its the same for every row Table 2 y/x = 3/1 = 3 6/2 = what?

OpenStudy (anonymous):

are you with me

OpenStudy (welshfella):

then check out 12/4 and 15/5 to these 4 come to the same value?

OpenStudy (welshfella):

do those simple divisions and you'll see

OpenStudy (fluttershyk):

ok im here

OpenStudy (anonymous):

okay good

OpenStudy (anonymous):

3*2= 3*4= 3*5=

OpenStudy (fluttershyk):

6 12 15

OpenStudy (welshfella):

thats it so what table is that?

OpenStudy (fluttershyk):

the second table

OpenStudy (welshfella):

oh i see you said the second one earlier Yes Table 2 is proportional and the proportionality constant is 3.

OpenStudy (fluttershyk):

ok thank u will you help with more

rebeccaxhawaii (rebeccaxhawaii):

got it?

OpenStudy (welshfella):

okz post more questions separately